Marigold Corp. can produce 100 units of a component part with the following costs:

Direct Materials $20000
Direct Labor 4500
Variable Overhead 18000
Fixed Overhead 11000


If Marigold Corp. can purchase the component part externally for $47000 and only $4000 of the fixed costs can be avoided, what is the correct make-or-buy decision?

a. Buy and save $2500

b. Make and save $500

c. Make and save $2500

d. Buy and save $4500
Answer #1

Cost of making = 20000+4500+18000 = 42500

Cost of buying = 47000-4000 = 43000

So, make and save 500 (42500-43000)

Option B is the answer